# Exact computation of Transfer Entropy with Path Weight Sampling

### v1

@@ Abstract
Information processing in networks entails a dynamical transfer of information between stochastic variables. Transfer entropy is widely used for quantification of the directional transfer of information between input and output trajectories. However, currently there is no exact technique to quantify transfer entropy given the dynamical model of a general network. Here we introduce an exact computational algorithm, Transfer Entropy-Path Weight Sampling (TE-PWS), to quantify transfer entropy and its variants in an arbitrary network in the presence of multiple hidden variables, nonlinearity, transient conditions, and feedback. TE-PWS extends a recently introduced algorithm Path Weight Sampling (PWS) and uses techniques from the statistical physics of polymers and trajectory sampling. We apply TE-PWS to linear and nonlinear systems to reveal how transfer entropy can overcome naive applications of data processing inequalities in presence of feedback.
@@
Information transfer between noisy signals underlies the functionality of diverse real-world networks such as in biochemical signaling, neuroscience, ecology, wireless communication and finance. Information theory has so far provided a useful framework to quantify the feasibility of retrieval of a noisy input signal from an output signal. In the presence of feedback loops in a network, information travels in both directions between the input and the output. An information-theoretic quantity that can measure the information transfer separately in either direction is the transfer entropy1, 2. Transfer entropy and its variants, such as directed information, conditional transfer entropy or filtered transfer entropy, have been widely used to gain knowledge about the connectivity of a network3, infer causal relations from experiments4, 5, establish fundamental bounds on network performance6, and estimate the minimal physical work required for a computation7, 8. Hence, for a wide range of problems, it is vital to be able to accurately quantify transfer entropies.
However, there are currently no exact methods to compute transfer entropy in a general many-variable network model beyond linear or low-order moment closure approximations9, 10. Systematic improvement of analytical approximations to account for high-order correlations in the trajectory ensemble is prohibitively complicated. Additionally, direct numerical estimation from experimental or simulated trajectories is intractable because the state space of long trajectories diverges exponentially with trajectory duration. Histograms in trajectory space are thus expensive to compute and are undersampled, giving rise to large and sometimes uncontrolled errors11, 12.
@@
Here we fill this gap by introducing TE-PWS, a numerical algorithm to estimate transfer entropies exactly in any stochastic model, including in diffusive and jump processes. The estimate is exact, i.e., it is an unbiased statistical estimate of the transfer entropy. Thus it can provide ground truth results for any given model. TE-PWS builds on the recently developed PWS algorithm for computation of mutual information between trajectories13. TE-PWS exploits the idea that path likelihoods can be obtained analytically from the Langevin or master equation, from which transfer entropy is then computed via Monte-Carlo averaging in trajectory space. Additionally, long trajectories are sampled with an importance sampling scheme, solving the problem of exponential computational cost. We have applied TE-PWS to compute the transfer entropy in a three-variable motif in the presence of feedback, for both linear and nonlinear dynamics. We show that TE-PWS reproduces analytical results when available and produces novel insights on how information feedback can amplify information transfer. Specifically, the transfer entropy from an input to an output node can overcome a naive application of data processing inequalities even when the mutual information obeys one.
@@
with $\mathbf{F}(t)$ a general drift, and $\bm{\xi}(t)$ a three-dimensional Gaussian white noise with a diffusion constant matrix $\mathbf{D}=[D_{ij}]$ such that $\langle\xi_{i}(t)\xi_{j}(t^{{}^{\prime}})\rangle=2D_{ij}\delta(t-t^{{}^{\prime}})$. The drift may depend on the entire past history as well as on time. The transfer entropy from $X_{i}$ to $X_{j}$ over $N$ timesteps of durations $\delta t$ each is defined as1
@@
where the index $k$ goes over individual timesteps, $X_{j}(k)$ denotes $X_{j}$ after $k$ timesteps, $X_{j,[0,k]}$ denotes the trajectory over the first $k$ timesteps, and $I(A;B)$ denotes the mutual information between two random variables $A$ and $B$. $\mathcal{T}_{X_{i}\to X_{j}}$ measures the information transferred from the past trajectory of $X_{i}$ to the new updates of $X_{j}$ at every timestep, given the past trajectory of $X_{j}$ is already known. In case the dynamics relaxes into a steady-state, we will also talk about the transfer entropy rate, $\dot{\mathcal{T}}_{X_{i}\to X_{j}}=\lim_{N\to\infty}\mathcal{T}_{X_{i}\to X_{i}}/(N\delta t)$.
@@
Figure 1: (a) Schematic representation of the increase in transfer entropy from variable $X_{i}$ to $X_{j}$ at the $(k+1)$-th timestep. Other variables $X_{l}$ may mediate information transfer even in the absence of a direct coupling from $X_{i}$ to $X_{j}$. (b) Propagation of reference dynamics for $X_{i}$ and $X_{l}$ such that it is commensurate with the given frozen $X_{j}^{(\nu)}$ trajectory. (c) In the RR scheme, trajectories are propagated from a reference distribution $P_{0}(X_{i,[0,k]},X_{l,[0,k]})$ and pruned and enriched periodically to turn them into the conditional distribution $P(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]}^{(\nu)})$.
For calculating transfer entropy using Eq. 4 we develop TE-PWS by extension of the PWS algorithm13. The central idea in TE-PWS is that trajectory likelihoods in the full $d$-dimensional space are analytically available on-the-fly, and that trajectory averages can be computed in a Monte-Carlo fashion. First, the average in Eq. 4 is computed as,
@@
where the index $\nu$ sums over $M_{1}$ pairs of trajectories of $X_{i}$ and $X_{j}$ sampled from the joint probability distribution $P\left(X_{i,[0,N]}^{(\nu)},X_{j,[0,N]}^{(\nu)}\right)$. For each pair of trajectories, the probabilties in the numerator and the denominator of Eq. 5 are not analytically available, but what is indeed available is the full joint probability $P(\mathbf{X}_{[0,N]}^{(\nu)})$ as the exponential of the Onsager-Machlup action14, 15. We thus need to marginalize over all degrees of freedom other than $X_{i}$ and $X_{j}$, denoted henceforth collectively as $X_{l}$. We illustrate this procedure first for the denominator in Eq. 5. It is obtained as $P(X^{(\nu)}_{j,[0,k+1]})/P(X^{(\nu)}_{j,[0,k]})$, where $P(X^{(\nu)}_{j,[0,k]})$ is obtained via marginalization,
@@
For performing this average in a Monte-Carlo fashion16, we sample from a reference distribution, $P_{0}(X_{i,[0,k]},X_{l,[0,k]})$, and correct the resultant bias by dividing by $P_{0}$,
@@
where the index $\mu$ sums over $M_{2}$ trajectories sampled from $P_{0}$. What is the best choice for $P_{0}$? The ideal choice would be the conditional distribution $P(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]})$, as it makes the summand in Eq. 7 equal for all $\mu$, such that the variance of the estimate of $P(X_{j,[0,k]}^{(\nu)})$ is zero. However, this conditional distribution is not known a priori. We therefore generate $X_{i}$ and $X_{l}$ trajectories in the frozen field of $X_{j,[0,N]}^{(\nu)}$ (Fig. 1b) resulting in a distribution $P_{0}$ that is known analytically and is expected to be close to this conditional distribution. To exactly compensate for the remaining deviations of $P_{0}$ from the conditional distribution, we employ, in the spirit of Rosenbluth-Rosenbluth(RR)-PWS13, a reweighing of the $X_{i}^{(\mu)}$ and $X_{l}^{(\mu)}$ trajectories on-the-fly with weights proportional to the ratio of the two distributions. We prune and enrich the trajectories with these weights after every $\delta t$ time. This gives us access to the conditional distribution exactly (see Appendix).
Returning now to the numerator in Eq. 5, it is an average over another conditional distribution,
@@
The first probability in the integral, which is the transition probability of $X_{j}$ in the full $d$-dimensional space, is analytically available. Additionally, similar to the procedure for the denominator in Eq. 5, $P(X_{l,[0,k]}|X_{i,[0,k]}^{(\nu)},X_{j,[0,k]}^{(\nu)})$ can also be accessed by sampling $X_{l}$ trajectories from a $P_{0}(X_{l,[0,k]})$, and applying the RR scheme. Thus the numerator in Eq. 5 can also be evaluated in a Monte-Carlo fashion (see Appendix).
@@
To summarize, for each of the $M_{1}$ pairs of $X_{i,[0,N]}$ and $X_{j,[0,N]}$ trajectories, we simulate a joint ensemble of $M_{2}$ new $X_{i}$ and $X_{l}$ trajectories to estimate the denominator in the logarithm in Eq. 5, and a separate ensemble of $M_{2}$ new $X_{l}$ trajectories to estimate the numerator (Fig. 1). The trajectories in each ensemble are pruned and enriched on-the-fly with the RR scheme, giving both the numerator and denominator in Eq. 5 as Monte-Carlo averages. The computational cost thus scales as $2M_{1}M_{2}$. The transfer entropy estimate is unbiased 13 and the statistical accuracy can be arbitrarily improved by increasing $M_{1}$ and $M_{2}$. A pseudocode for the algorithm is available in the Appendix. Aside from Schreiber’s transfer entropy1, other trajectory-based metrics of directional information transfer such as directed information17, conditional transfer entropy18 and filtered transfer entropy19, also can be derived from conditional distributions of trajectories. Hence TE-PWS can be used to compute all such metrics at similar cost, as shown in the Appendix.
@@
We demonstrate the validity of the method by analyzing two examples of an Ornstein-Uhlenbeck (OU) process for which transfer entropy rates are available. Eq. 1 is an OU process when it is linear with $\mathbf{F}=-\mathbf{a}\mathbf{X}$ where $\mathbf{a}$ is a spring constant matrix. The first example is a minimal model for describing the stochastic dynamics of gene expression and growth rate in bacteria20. The model is non-bipartite, meaning that the diffusion constant matrix is non-diagonal and the mutual information rates are not finite. Yet transfer entropy rates are finite and numerically exact results are available19. The other example is an OU process with two variables, for which transfer entropy rates are analytically available19. As shown in Fig. 2a and b, TE-PWS results converge to known transfer entropy rates within a few multiples of the relaxation timescales in the respective systems. Details about the models are provided in the Appendix.
@@
Figure 2: Convergence of transfer entropy rate estimates. (a) and (b) respectively show forward and reverse transfer entropy rates in the stochastic gene expression model in units of the inverse relaxation timescale $\mu_{0}$, as a function of the trajectory duration $T$. Symbols are from TE-PWS and dashed lines are numerically exact values. (c) and (d) show transfer entropy rates in the two-variable OU process. Symbols are from TE-PWS and dashed lines are analytical results.
@@
Jump process. For systems with jumps between a discrete number of states and finite waiting times between the jumps, such as a well-stirred chemical reaction network described by a master equation, or a neural spiking process, the dynamics is governed by a jump propensity matrix $\mathbf{Q}$ of dimensions $\mathcal{N}^{d}\times\mathcal{N}^{d}$, which describes jumps among the $\mathcal{N}$ states of each of the $d$ components of $\mathbf{X}$ 21. In contrast with diffusive processes which must be time-discretized to simulate, jump processes can be simulated exactly with event-driven kinetic Monte-Carlo algorithms such as the Gillespie algorithm22. Hence the transfer of information to $X_{j}$ during the $(k+1)$-th trajectory segment occurs at all instants of time within the segment $X_{j,[k,k+1]}$, rather than at only the endpoint $X_{j}(k+1)$. Therefore in the definition of the stepwise increments to transfer entropy, $\Delta\mathcal{T}_{[k,k+1]}^{i\to j}$, in Eq. 4, the logarithms of the probabilities $P(X_{j}(k+1)|X_{i,[0,k]},X_{j,[0,k]})$ and $P(X_{j}(k+1)|X_{j,[0,k]})$ should be replaced with functionals of the entire $X_{j,[k,k+1]}$ segment,
@@
where $\alpha$ counts the jumps that change the state of $X_{j}$, $\lambda_{ij}$ and $\lambda_{j}$ are escape propensities for $X_{j}$ in the marginal spaces of $(X_{i},X_{j})$ and $(X_{j})$ respectively, and $\mathcal{Q}_{ij}(\alpha)$ and $\mathcal{Q}_{j}(\alpha)$ similarly are marginal jump propensities for the $\alpha$-th jump2, 10. $Q_{ij},\lambda_{ij}$ and $Q_{j},\lambda_{j}$ are obtained by marginalizing jump propensities from the full $\mathbf{X}$-space into the $(X_{i},X_{j})$ and $(X_{j})$ spaces respectively2, 10.
@@
The transfer entropy in Eq. 11 can be calculated with TE-PWS by marginalizing jump propensities in a Monte-Carlo fashion over conditional distributions of hidden variables, similar to the diffusive case. The average is Eq. 11 is expressed as a Monte-Carlo average over simulated trajectories of $X_{i}$ and $X_{j}$, and each of $\mathcal{Q}_{ij},\lambda_{ij}$ and $\mathcal{Q}_{j},\lambda_{j}$ are expressed as Monte-Carlo averages over conditional distributions of hidden variables. Hidden variable trajectories are sampled from a $P_{0}$ propagated in the frozen field of the given trajectory, and the remaining bias is exactly corrected for through the RR scheme. Prior work argues that the escape terms involving $\lambda_{ij}$ and $\lambda_{j}$ in Eqs. 9 and 10 respectively can be omitted since they cancel each other on average2. We show however in the Appendix that the error in the transfer entropy estimate can be reduced by an order of magnitude by exploiting anti-correlated fluctuations between the escape and the jump terms.
@@
Data processing inequality. We demonstrate the utility of TE-PWS by applying it to a three-node motif to study whether transfer entropies obey Data Processing Inequalities (DPIs). Unidirectional flow of information between different nodes in a network leads to DPIs for mutual information23. For a general three-variable process, if the flow of information is $X_{1}\to X_{2}\to X_{3}$, i.e., without feedback, $I(X_{1,[0,N]};X_{3,[0,N]}|X_{2,[0,N]})=0$; here the right arrows denote flow of information as mediated either via activation or repression. This leads by the chain rule to $I(X_{1,[0,N]};X_{3,[0,N]})\leq I(X_{1,[0,N]};X_{2,[0,N]})$23. As transfer entropy equals mutual information in absence of feedback, it also obeys $\mathcal{T}_{X_{1}\to X_{3}}\leq\mathcal{T}_{X_{1}\to X_{2}}$. This relation bounds the amount of information that can be transmitted from input to output through an intermediate variable, yet is only valid in the absence of feedback. In presence of an $X_{2}\rightarrow X_{1}$ feedback, i.e., $X_{1}\rightleftarrows X_{2}\to X_{3}$, the mutual information continues to obey its DPI, while the transfer entropy formally doesn’t24, 25. This can be rationalized by considering the limit $X_{1}\leftarrow X_{2}\rightarrow X_{3}$, where $X_{2}$ controls both $X_{1}$ and $X_{3}$. Here the $X_{1}$ trajectories would still be predictive of fluctuations in $X_{3}$, thus $\mathcal{T}_{X_{1}\to X_{3}}\geq 0$ even as $\mathcal{T}_{X_{1}\to X_{2}}=0$. However, where the crossover from the feedforward to the feedback dominated regime occurs, and to what extent in practice transfer entropies can overcome the DPI, is currently not understood.
@@
Figure 3: (a) and (b) are typical $X_{1}$ (green) and $X_{2}$ (black) trajectories for models A and B at $a_{12}=2a_{21}=0.4$ and $a_{12}=a_{21}=-6$ respectively. (c) Transfer entropy rates in model A, denoted as $\dot{\mathcal{T}}_{X_{1}\to X_{3}}^{(\mathrm{A})}$ and $\dot{\mathcal{T}}_{X_{1}\to X_{2}}^{(\mathrm{A})}$, as a function of an increasing ratio of feedback to feedforward strength, $f^{*}$, when the feedforward strength is kept constant at $a_{21}=0.2$. (d) Similar as (c) but for model B, where the feedforward strength is kept constant at $a_{21}=-4$. (e) Ratio of transfer entropy rates overcomes DPI bound (dashed line at $\mathcal{T}^{*}=1$) with increasing feedback.
@@
We implement the motif $X_{1}\rightleftarrows X_{2}\to X_{3}$ with two diffusive models of mutual repression between $X_{1}$ and $X_{2}$, labeled A and B. In both models, $X_{3}$ rapidly copies the state of $X_{2}$ such that the information loss from $X_{2}$ to $X_{3}$ is low. Model A is a three-dimensional OU process with feedback, where the ratio of the feedback to feedforward spring constants $a_{12}/a_{21}=f^{*}$ is varied to study the violation of DPI. Model B is a nonlinear extension of A inspired by a genetic toggle switch26 where the drifts for $X_{1}$ and $X_{2}$ are changed to $F_{1}=-a_{11}X_{1}-a_{12}(1+X_{1}^{2})/(1+X_{1}^{2}+X_{2}^{2})$ and $F_{2}=-a_{22}X_{2}-a_{21}(1+X_{2}^{2})/(1+X_{1}^{2}+X_{2}^{2})$. We choose $a_{12}/a_{21}=f^{*}$ such that the two models can be compared. Other model parameters are provided in the Appendix. Typical $X_{1}$ and $X_{2}$ trajectories at high feedback are shown in Figs. 3a and b. Model A only shows regression to the mean for both $X_{1}$ and $X_{2}$, while model B additionally shows switching between (low,high) and (high,low) regimes for $X_{1}$ and $X_{2}$ for high values of both feedforward and feedback coupling, around $f^{*}=1$.

### v2

@@ Abstract
The ability to quantify the directional flow of information is vital to understanding natural systems and designing engineered information-processing systems. A widely used measure to quantify this information flow is the transfer entropy. However, until now, this quantity could only be obtained in dynamical models using approximations that are typically uncontrolled. Here we introduce a computational algorithm called Transfer Entropy-Path Weight Sampling (TE-PWS), which makes it possible, for the first time, to quantify the transfer entropy and its variants exactly for any stochastic model, including those with multiple hidden variables, nonlinearity, transient conditions, and feedback. By leveraging techniques from polymer and path sampling, TE-PWS efficiently computes the transfer entropy as a Monte-Carlo average over signal trajectory space. We apply TE-PWS to linear and nonlinear systems to reveal how transfer entropy can overcome naive applications of the data processing inequality in the presence of feedback.
@@
Information transfer between noisy signals underlies the functionality of diverse natural and man-made networks such as those in biochemical signaling, neuroscience, ecology, wireless communication and finance. Information theory has so far provided a useful framework for quantifying information transmission. In the presence of feedback loops in the network, information travels in both directions between the input and the output. An information-theoretic measure that can quantify the information transfer separately in either direction is the transfer entropy1, 2. Transfer entropy and its variants, such as directed information, conditional transfer entropy or filtered transfer entropy, have been widely used to gain knowledge about the connectivity of a network3, infer causal relations from experiments4, 5, establish fundamental bounds on network performance6, and estimate the minimal physical work required for a computation7, 8. Hence, for a wide range of problems, it is vital to be able to accurately quantify transfer entropies.
However, there are currently no exact methods to compute the transfer entropy in a general many-variable dynamical model. The transfer entropy depends on probability distributions of signal trajectories. Estimating these distributions by binning the experimental or simulated trajectories in histograms is not feasible, as the dimensionality of the trajectory space scales exponentially with signal duration 9, 10. As a result, different approximations are currently being used: either the full history dependence of the transfer entropy is truncated11, or an arbitrary distance metric in trajectory space is chosen for clustering the trajectories12, or a linear or low-order moment-closure approximation is employed13, 14, 15. These approximations can result in uncontrolled errors in complex, nonlinear, many-variable systems16, 17.
@@
Here we fill this gap by introducing TE-PWS, a numerical algorithm to estimate transfer entropies exactly for any stochastic model, including diffusive and jump processes. The estimate is exact, i.e., it is an unbiased statistical estimate of the transfer entropy. TE-PWS can therefore provide ground truth results for any given model. TE-PWS builds on the recently developed PWS algorithm for computing the mutual information between trajectories18. TE-PWS exploits the idea that path likelihoods can be obtained analytically from the Langevin or master equation, from which the transfer entropy is then computed via Monte-Carlo averaging in trajectory space. Additionally, long trajectories are sampled with an importance sampling scheme19, solving the problem of exponential scaling of the computational cost with trajectory duration. We have applied TE-PWS to compute the transfer entropy in a three-variable motif in the presence of feedback, for both linear and nonlinear systems. We show that TE-PWS reproduces analytical results when available and produces novel insights on how information feedback can amplify information transfer. Specifically, the transfer entropy from an input to an output node can overcome a naive application of the data processing inequality even when the mutual information obeys one.
@@
with $\mathbf{F}(t)$ a general drift, and $\bm{\xi}(t)$ a $d$-dimensional Gaussian white noise with a diffusion constant matrix $\mathbf{D}=[D_{ij}]$ such that $\langle\xi_{i}(t)\xi_{j}(t^{{}^{\prime}})\rangle=2D_{ij}\delta(t-t^{{}^{\prime}})$. The drift may depend on the entire past history as well as on time. The transfer entropy from $X_{i}$ to $X_{j}$ over $N$ timesteps of durations $\delta t$ each is defined as1
@@
where the index $k$ goes over individual timesteps, $X_{j}(k)$ denotes $X_{j}$ after $k$ timesteps, $X_{j,[0,k]}$ denotes the trajectory over the first $k$ timesteps, and $I(A;B)$ denotes the mutual information between two random variables $A$ and $B$. $\mathcal{T}_{X_{i}\to X_{j}}$ measures the information transferred from the past trajectory of $X_{i}$ to the new updates of $X_{j}$ at every timestep, given the past trajectory of $X_{j}$ is already known. In case the dynamics relaxes into a steady-state, we will also talk about the transfer entropy rate, $\dot{\mathcal{T}}_{X_{i}\to X_{j}}=\lim_{N\to\infty}\mathcal{T}_{X_{i}\to X_{j}}/(N\delta t)$.
@@
Figure 1: (a) Schematic representation of the increase in transfer entropy from variable $X_{i}$ to $X_{j}$ at the $(k+1)$-th timestep. Other variables $X_{l}$ may mediate information transfer even in the absence of a direct coupling from $X_{i}$ to $X_{j}$. (b) Propagation of reference dynamics for $X_{i}$ and $X_{l}$ such that it is commensurate with the given frozen $X_{j}^{(\nu)}$ trajectory. (c) In the RR scheme, trajectories sampled from a reference distribution $P_{0}(X_{i,[0,k]},X_{l,[0,k]})$ are resampled periodically to turn them into the desired conditional distribution $P(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]}^{(\nu)})$.
For calculating the transfer entropy using Eq. 4 we develop TE-PWS. Following PWS18, the central idea is that trajectory likelihoods in the full $d$-dimensional space are analytically available on-the-fly, and that trajectory averages can be computed in a Monte-Carlo fashion. First, the average in Eq. 4 is computed as,
@@
where the index $\nu$ sums over $M_{1}$ pairs of trajectories of $X_{i}$ and $X_{j}$ sampled from the joint probability distribution $P\left(X_{i,[0,N]}^{(\nu)},X_{j,[0,N]}^{(\nu)}\right)$. For each pair of trajectories, the probabilties in the numerator and the denominator of Eq. 5 are not analytically available, but what is indeed available is the full joint probability $P(\mathbf{X}_{[0,N]}^{(\nu)})$ as the exponential of the Onsager-Machlup action20, 21. We thus need to marginalize over all degrees of freedom other than $X_{i}$ and $X_{j}$, denoted henceforth collectively as $X_{l}$. We illustrate this procedure first for the denominator in Eq. 5. It is obtained as $P(X^{(\nu)}_{j,[0,k+1]})/P(X^{(\nu)}_{j,[0,k]})$, where $P(X^{(\nu)}_{j,[0,k]})$ is obtained via marginalization,
@@
For performing this average in a Monte-Carlo fashion19, we sample from a reference distribution, $P_{0}(X_{i,[0,k]},X_{l,[0,k]})$, and correct the resultant bias by dividing by $P_{0}$,
@@
where the index $\mu$ sums over $M_{2}$ trajectories sampled from $P_{0}$. What is the best choice for $P_{0}$? The ideal choice would be the conditional distribution $P(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]}^{(\nu)})$, as it makes the summand in Eq. 7 equal for all $\mu$, such that the variance of the estimate of $P(X_{j,[0,k]}^{(\nu)})$ is zero. However, this conditional distribution is not known a priori. We therefore generate $X_{i}$ and $X_{l}$ trajectories in the frozen field of $X_{j,[0,N]}^{(\nu)}$ (Fig. 1b) resulting in a distribution $P_{0}$ that is known analytically and is expected to be close to this conditional distribution. To exactly compensate for the remaining deviations of $P_{0}$ from the desired conditional distribution, we employ, in the spirit of Rosenbluth-Rosenbluth(RR)-PWS18, a reweighing of the $X_{i}^{(\mu)}$ and $X_{l}^{(\mu)}$ trajectories on-the-fly with weights proportional to the ratio of the two distributions. We resample the trajectories with these weights after every $\delta t$ time, meaning that we sample $M_{2}$ trajectories with replacement from the weighted trajectory ensemble (Fig. 1c). This procedure exactly generates $X_{i}^{(\mu)}$ and $X_{l}^{(\mu)}$ trajectories according to the desired conditional distribution $P(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]}^{(\nu)})$ (see End Matter (EM) and Supplemental Material (SM)).
Returning now to the numerator in Eq. 5, we note that it is an average over another conditional distribution,
@@
The first probability in the integral, which is the transition probability of $X_{j}$ in the full $d$-dimensional space, is analytically available. Additionally, similar to the procedure for the denominator in Eq. 5, samples from $P(X_{l,[0,k]}|X_{i,[0,k]}^{(\nu)},X_{j,[0,k]}^{(\nu)})$ are also available by sampling $X_{l}$ trajectories first from a $P_{0}(X_{l,[0,k]})$, and then applying the RR scheme. Thus the numerator in Eq. 5 can also be evaluated in a Monte-Carlo fashion (see EM).
@@
To summarize, for each of the $M_{1}$ pairs of $X_{i,[0,N]}$ and $X_{j,[0,N]}$ trajectories, we simulate a joint ensemble of $M_{2}$ new $X_{i}$ and $X_{l}$ trajectories to estimate the denominator in the logarithm in Eq. 5, and a separate ensemble of $M_{2}$ new $X_{l}$ trajectories to estimate the numerator (Fig. 1). The trajectories in each ensemble are resampled on-the-fly with the RR scheme, giving both the numerator and denominator in Eq. 5 as Monte-Carlo averages. The computational cost thus scales as $2M_{1}M_{2}$. The transfer entropy estimate is unbiased 18 and the statistical accuracy can be arbitrarily improved by increasing $M_{1}$ and $M_{2}$. A pseudocode for the algorithm is available in the SM. Aside from Schreiber’s transfer entropy1, other trajectory-based metrics of directional information transfer such as directed information22, conditional transfer entropy23 and filtered transfer entropy24, also can be derived from conditional distributions of trajectories. Hence TE-PWS can be used to compute all such metrics at similar cost, as shown in the SM.
@@
We demonstrate the validity of the method by analyzing two simple examples of an Ornstein-Uhlenbeck (OU) process for which transfer entropy rates are analytically available. Eq. 1 is an OU process when it is linear with $\mathbf{F}=-\mathbf{a}\mathbf{X}$ where $\mathbf{a}$ is a spring constant matrix. The first example is a minimal three-dimensional model for describing the stochastic dynamics of gene expression and the growth rate of bacteria25. The model is non-bipartite, meaning that the diffusion constant matrix is non-diagonal and the mutual information rates are not finite. Yet transfer entropy rates are finite and semi-analytically available24. The other example is an OU process with two variables, for which transfer entropy rates are analytically available (see EM). As shown in Fig. 2a and b, TE-PWS results converge to the known transfer entropy rates within a few multiples of the relaxation timescales in the respective systems. Details about the models are provided in the EM.
@@
Figure 2: Convergence of transfer entropy rate estimates. (a) and (b) respectively show forward and reverse transfer entropy rates in the stochastic gene expression model in units of an inverse relaxation timescale, $\mu_{0}^{-1}$, as a function of the trajectory duration $T$. Symbols are from TE-PWS and dashed lines are numerically exact values. (c) and (d) show transfer entropy rates in the two-variable OU process. Symbols are from TE-PWS and dashed lines are analytical results.
@@
Jump process. For systems with jumps between a discrete number of states and finite waiting times between the jumps, such as a well-stirred chemical reaction network described by a master equation, or a neural spiking process, the dynamics is governed by a jump propensity matrix $\mathbf{Q}$ of dimensions $\mathcal{N}^{d}\times\mathcal{N}^{d}$, which describes jumps among the $\mathcal{N}$ states of each of the $d$ components of $\mathbf{X}$ 26. In contrast with diffusive processes which must be time-discretized to simulate, jump processes can be simulated exactly with event-driven kinetic Monte-Carlo algorithms such as the Gillespie algorithm27. Hence the transfer of information from $X_{i}$ to $X_{j}$ during the $(k+1)$-th trajectory segment occurs at all instants of time within the segment $X_{j,[k,k+1]}$, rather than at only the endpoint $X_{j}(k+1)$. Therefore in the definition of the stepwise increments to transfer entropy, $\Delta\mathcal{T}_{[k,k+1]}^{i\to j}$, in Eq. 4, the logarithms of the probabilities $P(X_{j}(k+1)|X_{i,[0,k]},X_{j,[0,k]})$ and $P(X_{j}(k+1)|X_{j,[0,k]})$ should be replaced with functionals of the entire $X_{i,[k,k+1]}$ and $X_{j,[k,k+1]}$ segments,
@@
where $\alpha$ counts the jumps that change the state of $X_{j}$, $\lambda_{ij}$ and $\lambda_{j}$ are escape propensities for $X_{j}$ in the marginal spaces of $(X_{i},X_{j})$ and $(X_{j})$ respectively, and $\mathcal{Q}_{ij}(\alpha)$ and $\mathcal{Q}_{j}(\alpha)$ similarly are marginal jump propensities for the $\alpha$-th jump2, 15. $Q_{ij},\lambda_{ij}$ and $Q_{j},\lambda_{j}$ are obtained by marginalizing jump propensities from the full $\mathbf{X}$-space into the $(X_{i},X_{j})$ and $(X_{j})$ spaces respectively2, 15 (see SM).
@@
The transfer entropy in Eq. 11 can be calculated with TE-PWS by marginalizing jump propensities in a Monte-Carlo fashion over conditional distributions of hidden variables, similar to the diffusive case. Specifically, the average in Eq. 11 is expressed as a Monte-Carlo average over trajectories of $X_{i}$ and $X_{j}$, and each of $\mathcal{Q}_{ij},\lambda_{ij}$ and $\mathcal{Q}_{j},\lambda_{j}$ is expressed as a Monte-Carlo average over conditional distributions of hidden variables. Hidden variable trajectories are sampled from a $P_{0}$ propagated in the field of the frozen trajectory, and the remaining bias is exactly corrected for through the RR scheme (see SM). Prior work argues that the escape terms involving $\lambda_{ij}$ and $\lambda_{j}$ in Eqs. 9 and 10 respectively can be omitted since they cancel each other on average2. We show however in the SM that the error in the transfer entropy estimate can be reduced by an order of magnitude by exploiting anti-correlated fluctuations between the escape and the jump terms.
@@
Data processing inequality. We demonstrate the utility of TE-PWS by applying it to a three-node motif to study whether transfer entropies obey the Data Processing Inequality (DPI). Unidirectional flow of information between different nodes in a network leads to a DPI for the mutual information28. For a general three-variable process, if the flow of information is $X_{1}\to X_{2}\to X_{3}$, i.e., without feedback, $I(X_{1,[0,N]};X_{3,[0,N]}|X_{2,[0,N]})=0$; here the right arrows denote flow of information as mediated either via activation or repression. This leads by the chain rule to $I(X_{1,[0,N]};X_{3,[0,N]})\leq I(X_{1,[0,N]};X_{2,[0,N]})$28. As transfer entropy equals mutual information in the absence of feedback, it also obeys $\mathcal{T}_{X_{1}\to X_{3}}\leq\mathcal{T}_{X_{1}\to X_{2}}$. This relation bounds the amount of information that can be transmitted from input to output through an intermediate variable, yet is only valid in the absence of feedback. In the presence of an $X_{2}\rightarrow X_{1}$ feedback, i.e., $X_{1}\rightleftarrows X_{2}\to X_{3}$, the mutual information continues to obey its DPI, while the transfer entropy formally doesn’t29, 30. This can be rationalized by considering the limit $X_{1}\leftarrow X_{2}\rightarrow X_{3}$, where $X_{2}$ controls both $X_{1}$ and $X_{3}$. Here the $X_{1}$ trajectories would still be predictive of fluctuations in $X_{3}$, thus $\mathcal{T}_{X_{1}\to X_{3}}\geq 0$ even as $\mathcal{T}_{X_{1}\to X_{2}}=0$. However, where the crossover from the feedforward to the feedback-dominated regime occurs, and to what extent transfer entropies can overcome the DPI in practice, is currently not understood.
@@
Figure 3: (a) and (b) are typical $X_{1}$ (green) and $X_{2}$ (black) trajectories for models A and B at $a_{12}=2a_{21}=0.4$ and $a_{12}=a_{21}=-6$ respectively. (c) Transfer entropy rates in model A, denoted as $\dot{\mathcal{T}}_{X_{1}\to X_{3}}^{(\mathrm{A})}$ and $\dot{\mathcal{T}}_{X_{1}\to X_{2}}^{(\mathrm{A})}$, as a function of an increasing ratio of feedback to feedforward strength $f^{*}=a_{12}/a_{21}$, where the feedforward strength is kept constant at $a_{21}=0.2$. (d) Similar to (c) but for model B, where the feedforward strength is kept constant at $a_{21}=-4$. (e) Ratio of transfer entropy rates overcomes DPI bound (dashed line at $\mathcal{T}^{*}=1$) with increasing feedback.
@@
We implement the motif $X_{1}\rightleftarrows X_{2}\to X_{3}$ in two diffusive models of mutual repression between $X_{1}$ and $X_{2}$, labeled as models A and B. In both models, $X_{3}$ rapidly copies the state of $X_{2}$ such that the information loss from $X_{2}$ to $X_{3}$ is low. Model A is a three-dimensional OU process with linear feedback, where the ratio of the feedback to feedforward spring constants $f^{*}\equiv a_{12}/a_{21}$ is varied to study the violation of DPI. Model B is a nonlinear extension of A inspired by a genetic toggle switch31 where the drifts for $X_{1}$ and $X_{2}$ are changed to $F_{1}=-a_{11}X_{1}-a_{12}(1+X_{1}^{2})/(1+X_{1}^{2}+X_{2}^{2})$ and $F_{2}=-a_{22}X_{2}-a_{21}(1+X_{2}^{2})/(1+X_{1}^{2}+X_{2}^{2})$; we choose $a_{12}/a_{21}=f^{*}$ such that the two models can be compared. Other model parameters are provided in the EM. Typical $X_{1}$ and $X_{2}$ trajectories at high feedback are shown in Figs. 3a and b. Model A only shows regression to the mean for both $X_{1}$ and $X_{2}$, while model B additionally shows switching between (low,high) and (high,low) values of $(X_{1},X_{2})$ in the regime of both strong feedforward and feedback coupling, around $f^{*}=1$.